OpenStudy (automaticloveletter):

Consider the following pair of equations: −x − y = −5 y = x + 1 If the two equations are graphed, at what point do the lines representing the two equations intersect? (−2, 3) (3, −2) (2, 3) (3, 2)

OpenStudy (anonymous):

Use Geogebra.com or fill in answers. So -2-3=-5 and 3=-2+1 (NO -1 is not = to 3) IT IS NOT A 3+2=-5 no IT IS NOT B -2-3=-5 and 3=2+1 Those are both right the answer is C